If 100 mL of a .100 mol/L CaCl2(aq) and 100 mL of 0.0400 mol/L Na2SO4(aq) are mixed at 20°C, determine whether a precipitate will form. For CaSO4(aq) at 20°C, Ksp is 3.6 x 10-5.
Answer #1

Given data is :

volume of CaCl2 = 100 ml = 0.1 L

Molarity of CaCl2 = 0.100 mol/ L

Volume of Na2SO4 = 100 ml = 0.1L

Molarity of Na2SO4 = 0.0400 mol/L

Ksp for CaSO4 = 3.6 * 10^-5

Solution :

CaSO4 (aq ) ------> Ca2+(aq) + (So4)^2-(aq) at ksp = 3.6 * 10^-5

Moles of Ca2+ = molarity * volume in L

Moles of Ca2+ = 0.100 mol/L * 0.1L = 0.01 mol

Total volume = 0.1+ 0.1 = 0.2 L

Concentration of Ca2+ = moles / total volume

[Ca2+] = 0.01mol/ 0.2 L = 0.05mol/L

Moles of (So4)^2-= 0.0400 mol / L * 0.1L = 0.004 mol

Concentration of (So4)^2- = 0.004 mol/ 0.2 L =0.02 mol/L

Q = [Ca2+ ] [SO4^2-]

Q = 0.05 * 0.02 = 0.001 = 1 * 10^-3

Now

Q>ksp ; CaSO4 will precipitate from solution.